Consider observations of the number of wins per season by National Football League teams over the years 1989-2008. The observed mean is 8 wins/season and the observed standard deviation is 3.02. A histogram illustrates these data.

Referring to these NFL data and assuming normality, the probability that a randomly chosen team wins 10 or more games is?

Group of answer choices

a. 0.25

b. 0.16

c. 0.09

d. None of these options

e. 0.37
Answer #1

Solution :

Given that ,

mean = = 8

standard deviation = = 3.02

P(x > 10) = 1 - p( x< 10)

=1- p P[(x - ) / < (10 - 8) / 3.02 ]

=1- P(z < 0.66)

Using z table,

= 1 - 0.75

= 0.25

correct option is = a